OpenStudy (anonymous):

I need help geting started i did come up with anwser not correct, but i just need some instruction on how to begin the process. The forensic technician at a crime scene has just prepared a luminol stock solution by adding 17.0g of luminol into a total volume of 75.0 mL of H2O. What is the molarity of the stock solution of luminol? (Luminol has a molecular weight of 177 g/mol).

OpenStudy (aaronq):

use the molarity formula \(M=\dfrac{n_{solute}}{L_{solution}}\) \(n=\dfrac{m}{MW}\) so you can just plug one into the other to simplify things (or you can find moles first, then plug into the molarity formula). \(M=\dfrac{\dfrac{m}{MW}}{L_{solution}}=\dfrac{m}{MW*L_{solution}}\)

OpenStudy (anonymous):

i converted 75ml\[convert 75ml \to l = .075l h20, then convert 17.0 g luminol \to mol =.096mol luminol , then plug , the formula\] but i did something wrong should i do something with h20

OpenStudy (aaronq):

no thats fine, divide the moles by the liters now to find the molarity